### v1

# The distribution of interior branch lengths in phylogenetic trees under constant-rate birth-death models

Abstract
Phylogenetic trees represent the evolutionary relationships between extant lineages, omitting extinct or non-sampled lineages. Extending the work of Stadler and collaborators, this paper focuses on the distribution of branch lengths in phylogenetic trees arising under a constant-rate birth-death model. We derive branch length distributions of interior branches with and without random sampling of individuals of the extant population. We establish that branches connected to the tree leaves and interior branches behave very differently under sampling: pendant branches get longer as the sampling probability reaches zero, whereas the distribution of interior branches quickly reaches an asymptotic state.
I Introduction
The reconstructed evolutionary process is central to understanding evolutionary relationships and dynamics. This reconstruction is based on extant genetic data and yields a phylogenetic tree, which represents the evolutionary relationships among a set of species, genes, or cells. Crucially, all lineages that eventually went extinct (or were not sampled) are missing from this tree. This affects the statistics of the reconstructed tree: Under a stochastic birth-death process, species go extinct at a certain rate (or cells die, in a cellular system). Looking at the reconstructed tree, only speciation events are seen. To compare empirical phylogenetic trees with models of evolution, the link between an evolutionary process and the reconstructed process it induces must be understood.
The reconstructed evolutionary process has been studied extensively, with early work of Nee, May, and Harvey Nee et al. (1994) giving the probability density for each speciation event in a reconstructed tree from a birth-death process with constant rates at a particular time. Such a birth-death model describes a population of species, individuals, or cells, which divide independently at a fixed rate $\lambda$ and die independently at a fixed rate $\mu$. (For species, birth events describe the emergence of a new species, whereas death events correspond to extinctions.) A constant-rate birth-death model serves as a simple model of neutral evolution.
The work by Nee, May, and Harvey Nee et al. (1994) has been extended by Stadler and collaborators to trees reconstructed from individuals that were sampled from a population Stadler (2009) and to populations conditioned on a particular size Gernhard (2008). The statistics of speciation events have been used to infer trees and their characteristics Stadler (2013b); MacPherson et al. (2022), and is the cornerstone of widely-used computational packages like Beast Bouckaert et al. (2019) and TreeTime Sagulenko et al. (2018). Many extensions of the basic constant-rate birth-death process have been developed, for instance, to account for rates that are variable through time or differ across the tree (selection) Stadler (2013b).
In Stadler and Steel (2012); Mooers et al. (2012) Stadler, Steel and collaborators focus on a particular aspect of tree statistics, namely the distribution of branch lengths (in calendar time). They calculate the probability density function for the length of the so-called pendant branches (the lowest branches leading to the leaves of the tree) for trees reconstructed from a constant-rate birth-death process. For the particular case of a Yule process (a birth-death process with the rate of death set to zero), they also give the probability density function for the length of interior nodes.
In this paper, we revisit the problem of the distribution of branch lengths with a particular focus on the lengths of interior branches and the distribution of branch lengths under sampling of extant species. In Section II, we recapitulate the relevant properties of a constant-rate birth-death process and in Section III we introduce a simple but effective approach to calculate pendant branch lengths, recovering the results of Stadler and Steel Stadler and Steel (2012). We generalise these results to the lengths of interior branches: in Section IV branches one and two levels above a pendant branch are discussed, and in Section V we turn to generic branches picked randomly in the interior of a phylogenetic tree. In Section VI, we briefly discuss the statistics of the number of generations along branches, and in Section VII we generalise these results to a scenario where only a finite fraction of the extant population is sampled. We find that pendant and interior branch lengths behave very differently under sampling: while pendant branches become longer and grow without limit as the sampling probability approaches zero, the distribution of interior branch lengths quickly reaches an asymptotic limit as the sampling probability decreases.
II Dynamics of a birth-death process
We consider a birth-death process as introduced by Feller Feller (1939): individuals in a population duplicate with a constant rate of birth $\lambda$ and die with a constant rate of death $\mu$. The dynamics of the population size are described by a Markov process, with the probability $p_{n}(t)$ of the population consisting of $n$ individuals at time $t$ following a master equation
$\partial_{t}p_{n}(t)=\lambda(n-1)p_{n-1}(t)-\lambda np_{n}(t)+\mu(n+1)p_{n+1}(t)-\mu np_{n}(t)\ .$ (1)
Starting with a single individual at time $t=0$ and with a birth rate larger than the death rate $\lambda>\mu$, a population can grow exponentially with time, or it can die out at some time. Kendall Kendall (1948) derived the probability that such a population has died out by the time $t>0$
$p_{0}(t)=\mu\frac{1-\operatorname{\mathrm{e}}^{-(\lambda-\mu)t}}{\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)t}}\ ,$ (2)
as well as the probability that at time $t$, there is a single living individual
$p_{1}(t)=\frac{(\lambda-\mu)^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)t}}{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)t})^{2}}\ ,$ (3)
and the probability that at time $t$ the population has size $n$
$p_{n}(t)$ $=\left(1-p_{0}(t)\right)\left(1-\frac{\lambda}{\mu}p_{0}(t)\right)\left(\frac{\lambda}{\mu}p_{0}(t)\right)^{n-1}\ .$ (4)
The expected population size grows exponentially with
$\mathrm{E}\left[n(t)\right]=\sum_{n=0}^{\infty}np_{n}(t)=\operatorname{\mathrm{e}}^{(\lambda-\mu)t}\ .$ (5)
A powerful alternative to the master equation 1 is a master equation for the probability that a population that started with a single individual at time zero has population size one at time $t$ Maddison et al. (2007); Stadler (2009). Here, we generalise this approach to the probability that in a population that started with a single individual at time zero, at time $t$ there is one and only one individual that has surviving offspring at a later time $T$. This probability turns out to be a key tool to calculate the distribution of branch lengths.
Specifically, $\tilde{p}^{(T)}_{1}(t)$ denotes the probability that at time $t$ a clade which started at time zero contains one individual which has at least one extant descendant (including potentially itself) at time $T$, and all other individuals alive at time $t$ have no extant descendants at time $T$ (again including themselves). The master equation for $\tilde{p}^{(T)}_{1}(t)$ is
$\partial_{t}\tilde{p}^{(T)}_{1}(t)=-\mu\tilde{p}^{(T)}_{1}(t)-\lambda\left(1-2p_{0}(T-t)\right)\tilde{p}^{(T)}_{1}(t)\ ,$ (6)
where the first term accounts for the decay of probability due to the death of that individual, and the second term describes the probability that any additional lineage arising in a birth event fails to become extinct by the time of observation. The factor of two comes from the two offspring in such a birth event which can play this role. This is a master equation in a single probability only; the population aspect of this dynamics resides compactly in the probability $p_{0}(T-t)$ that a particular clade dies out in a time interval of length $T-t$ given by (2).
The master equation (6) is a first-order linear differential equation solved by the integrating factor $-\int_{0}^{t}\operatorname{\mathrm{d}\!}t^{\prime}\left[\mu+\lambda\left(1-2p_{0}(T-t^{\prime})\right)\right]$. Integrating to $t=T$ yields the probability of population size on at time $T$ (3). Integrating $t$ only to some intermediate time $T-\tau_{e}$ gives the probability $\tilde{p}^{(T)}_{1}(T-\tau_{e})$. (The name $\tau_{e}$ is chosen as this will later mark the end of an interior branch.)
It turns out to be convenient for the analysis of branch lengths to consider times relative to the time of observation $T$. We denote the length of the interval between the intermediate time $t$ and the observation time $T$ by $\tau_{e}=T-t$. We further define $\tau=T-\tau_{e}(=t)$ as the length of the interval between $\tau_{e}$ and the time of the clade’s origin. Specifically, we denote as $p_{1}^{(\tau_{e})}(\tau)=\tilde{p}^{(T)}_{1}(t=T-\tau_{e})$ the probability that a clade started at time $\tau_{e}+\tau$ in the past, and contains at time $\tau_{e}$ in the past only a single individual that will have extant descendants at the time of observation. Again, setting $\tau_{e}=0$ results in $p_{1}^{(\tau_{e}=0)}(\tau)\equiv p_{1}(\tau)$. For the case of $\tau_{e}>0$ the integrating factor becomes $-\int_{0}^{T-\tau_{e}}\operatorname{\mathrm{d}\!}t^{\prime}\left[\mu+\lambda\left(1-2p_{0}(T-t^{\prime})\right)\right]=-\int_{0}^{\tau}\operatorname{\mathrm{d}\!}\tau^{\prime}\left[\mu+\lambda\left(1-2p_{0}(\tau_{e}+\tau^{\prime})\right)\right]=\log\left(\frac{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}})^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{e}+\tau)})^{2}}\right)$ and we obtain
$p_{1}^{(\tau_{e})}(\tau)=\frac{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}})^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{e}+\tau)})^{2}}\ .$ (7)
This probability differs from $p_{1}(\tau)\equiv p^{(\tau_{e}=0)}_{1}(\tau)$, as for $\tau_{e}>0$ additional lineages have more time to die out than in pendant branches (specifically $\tau_{e}$ more). This will lead to a difference between the statistics of interior and pendant branches.
Figure 1: Times of events on a birth-death tree. Sketch of the situation characterized by $p_{1}^{(\tau_{e})}(\tau)=\tilde{p}_{1}^{(T)}(t)$. From the perspective of a clock starting at the origin of a clade, $p_{1}^{(\tau_{e})}(\tau)=\tilde{p}_{1}^{(T)}(t)$ describes the probability that out of all individuals at time $t$, only one will have extant descendants (depicted in blue) at time $T>t$ when the population is observed. The red lines show lineages which die out before $T$. For convenience, the time of observation is used as the origin of time $\tau$, with time increasing as we look into the evolutionary past. $\tau_{e}$ denotes time point beyond only a single individual has a clade extant at the time of observation, and $\tau$ denotes the time the population has had to grow before $\tau_{e}$. In the corresponding reconstructed phylogenetic tree, there are only the black and blue lineages.
In the following sections, we will always measure times $\tau$ relative to the time of observation, increasing as one looks further into the evolutionary past, see Figure 1
III Branch length distribution: pendant branches
The results from section II can be used directly to determine the statistics of branch lengths of a phylogenetic tree. In this section, we apply them to the so-called pendant branches, which end in a leaf of the phylogenetic tree. In this way, we recover the results of Stadler and Steel Stadler and Steel (2012). In the following sections, we apply the same technique to branches in the interior of a phylogenetic tree.
A pendant branch (a branch terminating at the time of observation) of length $\tau$ was created in a birth event at time $\tau$ , where (i) one of the two offspring produces a clade consisting of a single extant member at the time of observation, and (ii) the other offspring produces a clade that did not die out by the time of observation. Figure 2A shows a simple example. Conditions (i) and (ii) are statistically independent events whose probabilities can be expressed in terms of the quantities computed above. This leaves the rate at which birth events occur per small time interval, which is given by the birth rate $\lambda$ times the population size at a particular time.
We consider the case where the number of individuals $N$ at the time of observation is known, but no further information is given, for instance, on the age of the tree. Since the population size is unknown at time $\tau\neq 0$, we sum over all possible population sizes and their probabilities $\sum_{n=1}^{\infty}np_{n}(\tau)$. Ignatieva et al. (2020) have shown that the statistics of a population of size $N$ tracked backwards in time under a birth-death process is described by a birth-death process with time running forwards and birth and death rates exchanged. The expected population size of the forward process is proportional to $e^{(\lambda-\mu)t}$. Hence the expected population size at time $\tau$ before the observation is proportional to $\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}$. The dependence on $N$ (which can be summed over) is only in the constant of proportionality.
The relative number of pendant branches of length $\tau$ per small time interval is therefore given by
$\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}\left[p_{1}(\tau)p_{>1}(\tau)+p_{>1}(\tau)p_{1}(\tau)+2p_{1}(\tau)^{2}\right]$ (8)
$=2\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}p_{1}(\tau)\left(1-p_{0}(\tau)\right).$ (9)
The factor of two at the end of the first term accounts for the emergence of two branches of length $\tau$ when both offspring only have a single surviving descendant. $p_{>1}(\tau)=\sum_{n=2}^{\infty}p_{n}(\tau)$ denotes the probability that a clade founded at time $\tau$ has more than one extant member at the time of observation. Using equations (2) and (3), we obtain the probability density function of the lengths $\tau$ of pendant branches (up to a normalising factor found to be one) as
$P_{\text{pend}}(\tau)$ $=2\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}p_{1}(\tau)\left(1-p_{0}(\tau)\right)$ (10)
$=2\lambda\frac{(\lambda-\mu)^{3}\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}}{\left(\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}-\mu\right)^{3}}\ .$
Superficially, this result looks different from the distribution of pendant branch lengths derived in Stadler and Steel Stadler and Steel (2012) and Mooers et al. Mooers et al. (2012), which in the present notation is $2\lambda p_{1}(\tau)\left(1-\frac{\lambda}{\mu}p_{0}(\tau)\right)$. The results agree, as they must, since $\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}\left(1-p_{0}(\tau)\right)=\frac{(\lambda-\mu)\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}=1-\frac{\lambda}{\mu}p_{0}(\tau)$.
IV Branch length distribution: interior branches above pendant branches
This approach to pendant branches can be extended to the interior branches of a phylogenetic tree. We proceed to go up the phylogenetic tree and first consider branches one level up from pendant branches. We call branches that end in an internal node that is the beginning of a pendant branch level-two branches. Correspondingly, pendant branches are level-one branches. Figure 2B shows the situation: a pair of successive level-two and level-one branches of lengths $\tau_{2}$ and $\tau_{1}$, respectively, arose in a birth event at time $\tau_{2}+\tau_{1}$ in the past, where one of the two offspring has extant descendants (d-f), and the other has no extant descendants apart from the offspring of individuals arising in another birth event at time $\tau_{1}$, which gives rise to two offspring. Again, one of these has at least one extant descendant (b-c), and the other has exactly one (a), establishing the pendant branch of length $\tau_{1}$, see Fig. 2B.
ARefer to caption                    BRefer to caption
CRefer to caption
DRefer to caption      ERefer to caption
Figure 2: Statistics of pendant and interior branches. (A) This schematic phylogenetic tree shows birth events (circles) and death events ($\times$). The pendant branch of length $\tau$ on the left arose from a birth event in the past (solid circle), where one offspring (shown left) has exactly one extant descendant, and the other offspring (shown right) has at least one extant descendant. This particular pendant branch has one generation between its origin at time $\tau$ measured from the time of observation in the past and the present (not counting the birth event at $\tau$). (B) A birth event (top solid node) leading to a level-two branch of length $\tau_{2}$ and a level-one (pendant) branch of length $\tau_{1}$. (C) A schematic phylogenetic tree with branches of different levels, see text. (D) The distribution of branch lengths $\tau_{2}$ of second-level branches in a population of $N=7500$ individuals (blue bars), grown in a simulation as described in the text, compared to the theoretical probability density function (11) marginalised over the pendant branch lengths $\tau_{1}$ (red line). (E) The distribution of interior branch lengths in a population of $N=7500$ individuals (blue bars), compared to the theoretical density function (14) (red line).
Since the probability of a birth event at time $\tau_{1}+\tau_{2}$ was $\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{1}+\tau_{2})}$ (up to a normalising prefactor, see below), the relative number of level-two and level-one branches of lengths $\tau_{2}$ and $\tau_{1}$ per small time intervals is thus
$P_{\text{level}}(\tau_{1},\tau_{2})=2\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{1}+\tau_{2})}p_{1}(\tau_{1})\left(1-p_{0}(\tau_{1})\right)2\lambda p^{(\tau_{1})}_{1}(\tau_{2})\left(1-p_{0}(\tau_{1}+\tau_{2})\right)\ .$ (11)
Here $p^{(\tau_{1})}_{1}(\tau_{2})$ denotes the probability that all birth events taking place along the level-two branch of length $\tau_{2}$ lead to additional lineages that die out by the time the population is observed. More precisely, it is the probability that a given individual alive at time $\tau_{1}+\tau_{2}$ has exactly one descendant alive at time $\tau_{1}$ that will have extant descendants. It is given by equation (7).
Equation (11) gives the joint probability density function of branch lengths of pendant and level-two branches up to a normalisation factor, which again turns out to be one. Figure 2D compares this result to numerical simulations. A population with birth rate $\lambda=1$ and death rate $\mu=0.75$ was grown from a single individual to a population of $7500$ individuals (stopping when the population first reaches that number), and the phylogenetic tree was reconstructed. The lengths of level-two branches were collected in a histogram and compared to the joint probability $P_{\text{level}}(\tau_{1},\tau_{2})$ given by (11) marginalised with respect to $\tau_{1}$. We note that this simulation produces statistics that are slightly different from the statistics derived above. The simulation halts when a particular population size is reached, whereas the calculation looks at the statistics at a particular time $T$, see Hartmann et al. (2010) for the subtle distinction. At large population sizes, the fluctuations around the mean population size and, hence, the difference between these two ensembles are small.
This approach can be applied to higher levels in the phylogenetic tree. The joint probability density function of branch lengths $\tau_{1},\tau_{2}$ and $\tau_{3}$ in a lineage with level-one, -two, and -three branch is
$P_{\text{level}}(\tau_{1},\tau_{2},\tau_{3})=2\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{1}+\tau_{2}+\tau_{3})}p_{1}(\tau_{1})\left(1-p_{0}(\tau_{1})\right)2\lambda p^{(\tau_{1})}_{1}(\tau_{2})\left(1-p_{0}(\tau_{1}+\tau_{2})\right)2\lambda p^{(\tau_{1}+\tau_{2})}_{1}(\tau_{3})\left(1-p_{0}(\tau_{1}+\tau_{2}+\tau_{3})\right)\ .$ (12)
These densities describe the statistics of branches as one goes up the tree from the different leaves; starting from one particular leaf gives one set of branch lengths $\tau_{1},\tau_{2},\ldots$, and starting from another leaf gives another set of branch lengths. However, at higher levels than the first, not all branches can be uniquely assigned to a specific level. For instance, the branch marked $\ast$ in Figure 2C contributes both to the distribution of $\tau_{2}$ (going up the tree from the leaf $a$) and to the distribution of $\tau_{3}$ (going up from the leaves marked $b$ or $c$).
V Branch length distribution: interior branches
As a last step, we derive the distribution of internal branch lengths, i.e. of all branches that are not pendant. To this end, we calculate the joint probability of an interior branch starting at time $\tau_{s}$ and ending at $\tau_{e}<\tau_{s}$. As before, the probability per unit time of a birth event occurring at time $\tau_{s}$ is proportional to $\lambda\operatorname{\mathrm{e}}^{-(1-q)\tau_{s}}$. Focusing for now on only one of the two clades that started at $\tau_{s}$ (called $A$), the probability of it running until $\tau_{e}$ with only one extant descendant at the time of observation is $p_{1}^{(\tau_{e})}(\tau_{s}-\tau_{e})$. The probability per small time interval of the branch ending at $\tau_{e}$ is equal to the probability per small time interval of a birth event at that time where none of the two emerging clades dies out. This is given by $\lambda(1-p_{0}(\tau_{e}))^{2}$.
Two cases have to be considered for the other clade (called $B$) emerging at $\tau_{s}$. With probability $p_{1}(\tau_{s})$ this clade becomes a pendant branch. In this case, a factor of two arises due to symmetry (either clade can be called $A$). An alternative scenario is the clade having more than one extant descendant. This happens with probability $p_{>1}(\tau_{s})=1-p_{0}(\tau_{s})-p_{1}(\tau_{s})$. In this case, the clade also has an interior branch starting at $\tau_{s}$ and therefore contributes to the branch length distribution. A factor of two accounts for this.
Putting these steps together, the relative probability of an interior branch running from $\tau_{s}$ to $\tau_{e}$ (up to a normalising factor found to be one) is given by
$P_{\text{int}}(\tau_{s},\tau_{e})$ $=\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{s}}p_{1}^{(\tau_{e})}(\tau_{s}-\tau_{e})\lambda\left(1-p_{0}(\tau_{e})\right)^{2}\left[2p_{1}(\tau_{s})+2p_{>1}(\tau_{s})\right]$
$=2\lambda^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{s}}p_{1}^{(\tau_{e})}(\tau_{s}-\tau_{e})\left(1-p_{0}(\tau_{e})\right)^{2}\left(1-p_{0}(\tau_{s})\right)$ (13)
$=2\lambda^{2}(\lambda-\mu)^{3}\frac{\operatorname{\mathrm{e}}^{(\lambda-\mu)(\tau_{s}+\tau_{e})}}{(\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau_{s}}-\mu)^{3}}\ .$
Marginalizing over $\tau_{s}$ for $\tau=\tau_{s}-\tau_{e}$ results in the density of interior branch lengths in the entire tree
$P_{\text{int}}(\tau)$ $=(\lambda-\mu)^{2}\frac{2\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{\left(\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}-\mu\right)^{2}}.$ (14)
Figure 2E compares this result to the empirical distribution of a tree reconstructed from a population of $7500$ individuals grown from a single individual with birth rate $\lambda=1$ and death rate $\mu=0.75$ as described above.
VI Number of generations along branches
Birth events where one of the two newly founded clades dies out do not lead to bifurcations in the reconstructed tree. Such birth events are not visible from the reconstructed tree alone, yet they contribute to the number of generations along a lineage. In this section, we discuss the statistics of the number of generations along a particular branch of a phylogenetic tree.
The mean number $i$ of generations along a particular branch starting at time $\tau_{s}$ and ending at time $\tau_{e}$ (not counting the founding event of the branch) has been calculated by Bokma, van den Brink, and Stadler Bokma et al. (2012) as
$\langle i\rangle(\tau_{s},\tau_{e})=2\lambda\int_{\tau_{e}}^{\tau_{s}}\operatorname{\mathrm{d}\!}\tau\,p_{0}(\tau)=2\left[\lambda\tau-\log\left(\lambda-\mu e^{-(\lambda-\mu)\tau}\right)\right]_{\tau_{e}}^{\tau_{s}}$ (15)
based on birth events occurring along a branch at rate $\lambda$ and the probability $p_{0}(\tau)$ that one given additional lineage has died out by time $\tau$ given by (2). The factor $2$ arises because either of the two offspring can be the one to die out. The number of generations is Poisson distributed (and hence fully characterised by the mean (15)), because successive birth events are statistically independent Bokma et al. (2012).
Together with the densities of branch lengths from the previous sections, we can now compute the probability distribution for the number of generations along a randomly chosen branch. In the case of pendant branches, it is given by
$P_{\text{pend}}(i)=\int_{0}^{\infty}\operatorname{\mathrm{d}\!}\tau\,P_{\text{pend}}(\tau)\operatorname{\mathrm{e}}^{-\langle i\rangle(\tau,0)}\frac{\left[\langle i\rangle(\tau,0)\right]^{i}}{i!}\ .$ (16)
Similarly, the probability of finding $i$ generations along a randomly chosen interior branch is
$P_{\text{int}}(i)=\int_{0}^{\infty}\operatorname{\mathrm{d}\!}\tau_{s}\int_{0}^{\tau_{s}}\operatorname{\mathrm{d}\!}\tau_{e}\,P_{\text{int}}(\tau_{s},\tau_{e})\operatorname{\mathrm{e}}^{-\langle i\rangle(\tau_{s},\tau_{e})}\frac{\left[\langle i\rangle(\tau_{s},\tau_{e})\right]^{i}}{i!}\ .$ (17)
VII Finite sampling
In many practical applications, the phylogenetic tree is not based on the entire extant population at the time of observation but on a subset of individuals that were sampled at that time. Reconstructing the phylogenetic tree from the sampled individuals changes the statistics of branch lengths derived above. A standard model for the process of sampling from the extant population is to choose extant individuals at the time of observation independently and with equal probability $\rho$ Yang and Rannala (1997); Stadler (2010); Stadler and Steel (2012). The preceding results on branch length distributions can be easily generalised to a sampling probability $\rho<1$ using the probabilities that a clade of a particular age has no (or one) member which is sampled at a particular time.
Yang and Rannala Yang and Rannala (1997) have calculated the probabilities that a birth-death process starting with a single individual at time zero has no sampled offspring or a single sampled offspring, respectively, as
$p_{0}^{(\rho)}(\tau)$ $=1-\frac{\rho(\lambda-\mu)}{\rho\lambda+\left((1-\rho)\lambda-\mu\right)\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}$ (18)
$p_{1}^{(\rho)}(\tau)$ $=\frac{\rho(\lambda-\mu)^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{(\rho\lambda+\left((1-\rho)\lambda-\mu\right)\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau})^{2}}\ .$ (19)
[…]
$p_{1}^{(\rho)(\tau_{e})}(\tau)$ $=\frac{\left(\rho\lambda-(\mu-(1-\rho)\lambda)\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}}\right)^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{\left(\rho\lambda-(\mu-(1-\rho)\lambda)\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{e}+\tau)}\right)^{2}}\ .$ (20)
To derive the probability density function of branch lengths under sampling, these probabilities can be used directly in the densities of pendant and interior branch lengths at full sampling ($\rho=1$), (10)-(13), which depend only on $p_{0}$, $p_{1}$, and the relative population sizes at different times. In particular, we get
$P^{(\rho)}_{\text{pend}}(\tau)$ $=2\lambda\rho(\lambda-\mu)^{3}\frac{\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}}{\left(\lambda\rho\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}+\lambda(1-\rho)-\mu\right)^{3}}$ (21)
$P_{\text{int}}^{(\rho)}(\tau_{s},\tau_{e})$ $=2(\lambda\rho)^{2}(\lambda-\mu)^{3}\frac{\operatorname{\mathrm{e}}^{(\lambda-\mu)(\tau_{s}+\tau_{e})}}{\left(\lambda\rho\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau_{s}}+\lambda(1-\rho)-\mu\right)^{3}}$ (22)
$P_{\text{int}}^{(\rho)}(\tau)$ $=(\lambda-\mu)^{2}\frac{2\lambda\rho-\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}\left(\mu-\lambda(1-\rho)\right)}{\left(\lambda(1-\rho)-\mu+\lambda\rho\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}\right)^{2}}$ (23)
Note that a normalising factor of $\rho^{-1}$ applies to both densities of interior branches, while the normalising factor in the case of pendant branches again turns out to be one. The case of pendant branches extends the result of Stadler and Steel Stadler and Steel (2012), which was restricted to a regime of high death rate or high sampling probability $1\geq\mu/\lambda\geq 1-\rho$. The density function (22) gives the same result but also holds outside these limits.
Stadler Stadler (2009) found that the probability distribution of bifurcation times is not unique to every set of parameters, but that a birth-death process with birth rate $\lambda$, death rate $\mu$ and sampling probability $\rho$ yields the same distribution as a process with parameters $\rho^{\prime}$, $\lambda^{\prime}=\rho\lambda/\rho^{\prime}$, $\mu^{\prime}=\mu-\lambda(1-\rho/\rho^{\prime})$. We find the same for the distributions of branch lengths (21)-(23), as $\lambda^{\prime}-\mu^{\prime}=\lambda-\mu$, $\rho^{\prime}\lambda^{\prime}=\rho\lambda$ and $\lambda^{\prime}(1-\rho^{\prime})-\mu^{\prime}=\lambda(1-\rho)-\mu$ (which is the same argument as in Stadler (2009, 2013a)). The inference of both growth parameters and the sampling probability from the distribution of branch lengths is therefore not possible. However, the maximal sampling probability can be estimated: Given a branch distribution generated with (unknown) parameters $(\lambda,\mu,\rho)$, the parameter space of all sets of $(\lambda^{\prime},\mu^{\prime},\rho^{\prime})$ which lead to the same distribution is restricted to $\mu^{\prime}\geq 0$ and $0<\rho\leq 1$. With the transformation rules above, it is easy to see that these inequalities are only fulfilled if $0<\rho^{\prime}\leq\min(\rho\lambda/(\lambda-\mu),1)$. This means that for branch distributions generated with birth rate $\lambda$, death rate $\mu$ and sampling probability $\rho<1-\mu/\lambda$, there exists a maximal sampling probability $\rho^{\prime}<1$ for the family of alternative parameter sets.
The lengths of pendant and interior branches behave very differently under sampling. Both cases are depicted in Figure 3A. The rows correspond to different sampling probabilities $\rho$. Histograms show the branch length distributions of 100 simulations of trees with 300 sampled individuals, while the orange dashed lines depict the analytic results. Figure 3A shows in the second column the length distribution of pendant branches, which get longer with decreasing sampling probability $\rho$. On the other hand, interior branches quickly reach an asymptotic distribution as $\rho$ is decreased and then do not get any longer as $\rho$ is decreased further, see the first column of Figure 3A. The red dotted line depicts the asymptotic state, which can be derived as
$P^{(\rho\rightarrow 0)}_{\text{int}}(\tau)=\lim_{\rho\rightarrow 0}P^{(\rho)}_{\text{int}}(\tau)=(\lambda-\mu)\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}$ (24)
and is thus an exponential distribution with parameter $\lambda-\mu$.
ARefer to caption BRefer to caption CRefer to caption
Figure 3: Pendant and interior branches at different sampling probabilities $\rho$. (A) We compare the analytical results for the density of the length of interior and pendant branches under sampling, (23) and (21) respectively, to numerical simulations. For the latter, we show the distribution of branches from 100 trees, each with 300 pendant branches. For every tree, we simulated the growth of a population with $\lambda=1$ and $\mu=0.75$ up to a size of $300/\rho$ and reconstructed the tree based on 300 randomly chosen individuals. (Fixing both population and sample size in this way is distinct from independent sampling of each individual. However, the difference vanishes with increasing population size.) Trees are grown in this way for different values of the sampling probability $\rho$, depicted in the different rows. The first column shows the distribution of interior branches, quickly reaching the asymptotic distribution (24) (red dotted line) as $\rho$ becomes small. The second column shows the distributions of pendant branch lengths, which shift to longer branches as $\rho$ decreases. (B) The distribution of the number of generations along interior branches in the asymptotic case of small sampling probabilities (25). Different colours represent different relative death rates $\mu/\lambda$. (C) A toy example of how sampling (here at sampling probability $\rho=1/2$) affects the length of pendant branches, see text.
Following the steps in Section VI using equations (21)-(23), we can compute the distribution of generations along branches of a sampled tree. Crucially, since the mean number of generations (equation (15)) along a branch of length $\tau$ converges to $2\mu\tau$ as the branch moves further away from the time of observation (as interior branches do when taking the limit $\rho\rightarrow 0$), we can compute the probability of finding $i$ generations along an interior branch in this limit
$P^{(\rho\rightarrow 0)}_{\text{int}}(i)=\int_{0}^{\infty}\operatorname{\mathrm{d}\!}\tau\,P^{(\rho\rightarrow 0)}_{\text{int}}(\tau)\operatorname{\mathrm{e}}^{-2\mu\tau}\frac{(2\mu\tau)^{i}}{i!}=(2\mu)^{i}(\lambda-\mu)(\lambda+\mu)^{-i-1}\ .$ (25)
This result is depicted for different relative death rates $\mu/\lambda$ in Figure 3B.
The difference between pendant and interior branches under sampling has a simple qualitative explanation, which we explore within a toy problem. Consider a reconstructed tree which is perfectly balanced; both clades defined by each bifurcation have the same size. An example is shown in Fig. 3C. Then a fraction $\rho$ of the extant nodes is sampled. Under sampling, a pair of pendant branches linked to a shared ancestor is defined by the birth of a clade (shown in blue in Fig. 3C) whose size $n$ at the time of observation is defined by $n\rho\approx 2$ (so two of its members are expected to be sampled). The number of branch levels between this founding birth event and the time of observation is $\log_{2}(n)$, so the pendant branch lengths are increased under sampling by a factor of $1-\log_{2}\rho\approx-\log_{2}\rho$ relative to the fully sampled case. (This argument assumes that on the fully sampled reconstructed tree, pendant and interior branches are approximately of the same length.) On the other hand, branches that are interior in the reconstructed tree (i.e. lie above birth event shown in blue in Fig. 3C) have their lengths unchanged by sampling. The distribution of interior branches only changes because under sampling, fewer and older branches are counted as interior. Older branches tend to be a little longer (see equation (13)), because any additional lineages that arose along them had more time to die out until the time of observation.
VIII Discussion
We have calculated the statistics of branch lengths (in calendar time) of phylogenetic trees generated by a constant-rate birth-death process and reconstructed from the population at a single moment in time. Our approach simplifies the derivation of the distribution of pendant branch lengths by Stadler and Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012) and allows to treat both branches at a given distance from the tree leaves as well as the branch length distribution of randomly chosen interior branches. We further generalised the distributions to the case of finite sampling. While we derived our distributions conditioned on the population size at the time of sampling, they turn out to be independent of that size. Other cases, such as conditioning on the first bifurcation time or on both bifurcation time and population size are still open for further investigation.
We found that pendant branches and interior branches behave very differently under sampling: only pendant branch lengths increase indefinitely as the sampling probability is decreased to zero, whereas the distribution of interior branches quickly reaches an asymptotic. This behaviour allows to place bounds on the model parameters (birth and death rates, sampling probability) on the basis of empirical trees.
The behaviour of interior branch lengths also has consequences for the number of generations along the branches: even at low sampling probabilities, the number of generations along interior branches can be small — depending on the ratio of death and birth rates, but asymptotically independently of the sampling. Under cellular reproduction, when mutations occur at birth events, this means that the number of mutations along a branch can be shaped by the (few) birth events along that branch rather than accumulating continuously over time. The consequences and applications of this result will be explored in a future publication.
Acknowledgements.
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) grant SFB1310/2 - 325931972. Many thanks to Tibor Antal, Samuel Johnston, Joachim Krug for discussions, and to Tanja Stadler for comments throughout and especially for helpful advice on section VII. We thank the computing center of the University of Cologne RRZK for computing time on the DFG-funded CHEOPS cluster (grant number INST 216/512/1FUGG).