### v2

# Branch length statistics in phylogenetic trees under constant-rate birth-death models

Abstract
Phylogenetic trees represent the evolutionary relationships between extant lineages, where extinct or non-sampled lineages are omitted. Extending the work of Stadler and collaborators, this paper focuses on the branch lengths in phylogenetic trees arising under a constant-rate birth-death model. We derive branch length distributions of phylogenetic branches with and without random sampling of individuals of the extant population under two distinct statistical scenarios: a fixed age of the birth-death process and a fixed number of individuals at the time of observation. We find that branches connected to the tree leaves (pendant branches) and branches in the interior of the tree behave very differently under sampling: pendant branches get longer as the sampling probability is decreased, whereas the distribution of interior branches quickly reaches an asymptotic regime.
I Introduction
The reconstructed evolutionary process is central to understanding evolutionary relationships and dynamics. This reconstruction is based on extant genetic data and yields a phylogenetic tree, which represents the evolutionary relationships among a set of species, genes, or cells. Crucially, all lineages that eventually went extinct (or were not sampled) are missing from such a tree. This affects the statistics of the reconstructed tree: Under a stochastic birth-death process, species go extinct at a specific rate (or cells die, in a cellular system). Looking at the reconstructed tree, however, only speciation events are seen. To compare empirical phylogenetic trees with models of evolution, it is crucial to understand the link between an evolutionary process and the reconstructed process.
The reconstructed evolutionary process has been studied extensively, with early work of Nee, May, and Harvey Nee et al. (1997) giving the probability density in time for each speciation event in a reconstructed tree from a birth-death process with constant rates. Such a birth-death model describes a population of species, individuals, or cells, which divide independently at a fixed rate $\lambda$ and die independently at a fixed rate $\mu$. (For species, birth events describe the emergence of a new species, whereas death events correspond to extinctions.) A constant-rate birth-death model serves as a simple model of neutral evolution.
Stadler and collaborators have extended the work by Nee, May, and Harvey Nee et al. (1997) to trees reconstructed from individuals that were sampled from a population Stadler (2009) and to populations conditioned on a particular size Gernhard (2008). The resulting statistics of speciation events have been used to infer trees and their characteristics Stadler (2013b); MacPherson et al. (2022), and is the cornerstone of widely-used computational packages such as BEAST Bouckaert et al. (2019) and TreeTime Sagulenko et al. (2018). Many extensions of the basic constant-rate birth-death process have been developed, for instance, to account for rates that are variable through time or vary across the tree (selection) Stadler (2013b).
In Stadler and Steel (2012); Mooers et al. (2012) Stadler, Steel and collaborators focus on a particular aspect of tree statistics, namely the distribution of branch lengths (in calendar time). They calculate the probability density function for the length of the so-called pendant branches (the lowest branches leading to the leaves of the tree) for trees reconstructed from a constant-rate birth-death process. For the particular case of a Yule process (a birth-death process with the rate of death set to zero), they also give the probability density function for the length of interior branches.
In this paper, we revisit the problem of the distribution of branch lengths with a particular focus on the lengths of interior branches and the distribution of branch lengths under sampling of extant species. Two statistical scenarios need to be distinguished: In the first scenario (scenario i)), the time interval between the start of the birth-death process, starting with a single individual, and the time of observation (the end of the birth-death process) is known, but the number of extant species at the end of that process is a fluctuating random variable. In the second scenario (scenario ii)), one conditions on the number of individuals at the end of the process, but the time interval between the start and end of the process is unknown.
In Section II, we recapitulate the relevant properties of a constant-rate birth-death process and in Sections III-V, we discuss a simple but effective approach to determine branch length distributions under the first scenario used by Paradis Paradis (2016). We write down concrete expressions for branch length distributions in the limit of a long-running birth-death process, focussing on pendant branches in Section III, on branches one and two levels above a pendant branch in Section IV, and on generic branches picked randomly in the interior of a phylogenetic tree in Section V. Section VI briefly discusses the statistics of the number of generations along branches. In Section VII we generalize these results to a scenario where only a finite fraction of the extant population is sampled. We find that pendant and interior branch lengths behave very differently under sampling: while pendant branches become longer and grow without limit as the sampling probability approaches zero, the distribution of interior branch lengths quickly reaches an asymptotic limit as the sampling probability decreases. In Section VIII we turn to the second scenario (conditioning on the number of extant species at the time of observation) and introduce a novel approach based on tracking the birth-death process first backwards in time from the time of observation to the emergence of a particular branch and then again forwards to the time of observation. For pendant branches, we recover the results of Stadler, Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012) and generalize these results to interior branches.
II Dynamics of a birth-death process
We consider a birth-death process as introduced by Feller Feller (1939): individuals in a population duplicate with a constant rate of birth $\lambda$ and die with a constant rate of death $\mu$. The dynamics of the population size are described by a Markov process, with the probability $p_{n}(t)$ of the population consisting of $n$ individuals at time $t$ following a master equation
$\partial_{t}p_{n}(t)=\lambda(n-1)p_{n-1}(t)-\lambda np_{n}(t)+\mu(n+1)p_{n+1}(t)-\mu np_{n}(t)\ .$ (1)
Starting with a single individual at time $t=0$ and with a birth rate larger than the death rate $\lambda>\mu$, a population can grow exponentially with time, or it can die out at some time. Kendall Kendall (1948) derived the probability that such a population has died out by the time $t>0$
$p_{0}(t)=\mu\frac{1-\operatorname{\mathrm{e}}^{-(\lambda-\mu)t}}{\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)t}}\ ,$ (2)
as well as the probability that at time $t$, there is a single living individual
$p_{1}(t)=\frac{(\lambda-\mu)^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)t}}{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)t})^{2}}\ ,$ (3)
and the probability that at time $t$ the population has size $n$
$p_{n}(t)$ $=\left(1-p_{0}(t)\right)\left(1-\frac{\lambda}{\mu}p_{0}(t)\right)\left(\frac{\lambda}{\mu}p_{0}(t)\right)^{n-1}\ .$ (4)
The expected population size grows exponentially with
$\mathrm{E}\left[n(t)\right]=\sum_{n=0}^{\infty}np_{n}(t)=\operatorname{\mathrm{e}}^{(\lambda-\mu)t}\ .$ (5)
A powerful alternative to the master equation (1) is a master equation for the probability that a population that started with a single individual at time zero has population size one at time $t$ Maddison et al. (2007); Stadler (2010). A generalization of this approach is the probability that in a population that started with a single individual at time zero, at time $t$ there is one and only one individual that has surviving offspring at a later time $T$. This probability turns out to be a key tool to calculate the distribution of branch lengths.
Specifically, $\tilde{p}^{(T)}_{1}(t)$ denotes the probability that at time $t$ a clade which started at time zero contains one individual with open future fate, and all other individuals alive at time $t$ have no extant descendants at time $T$ (including themselves). The master equation for $\tilde{p}^{(T)}_{1}(t)$ is
$\partial_{t}\tilde{p}^{(T)}_{1}(t)=-\mu\tilde{p}^{(T)}_{1}(t)-\lambda\left(1-2p_{0}(T-t)\right)\tilde{p}^{(T)}_{1}(t)\ ,$ (6)
where the first term accounts for the decay of probability due to that individual’s death, and the second term describes the probability that any additional lineage arising in a birth event fails to become extinct by the time of observation. The factor of two comes from the two offspring in such a birth event which can play this role. This is a master equation in a single probability only; the population aspect of this dynamics resides compactly in the probability $p_{0}(T-t)$ that a particular clade dies out in a time interval of length $T-t$ given by (2).
The master equation (6) is a first-order linear differential equation solved by the integrating factor $-\int_{0}^{t}\operatorname{\mathrm{d}\!}t^{\prime}\left[\mu+\lambda\left(1-2p_{0}(T-t^{\prime})\right)\right]$. Integrating to $t=T$ yields the probability of population size one at time $T$ (3). Integrating $t$ only to some intermediate time $T-\tau_{e}$ gives the probability $\tilde{p}^{(T)}_{1}(T-\tau_{e})$. (The name $\tau_{e}$ is chosen as this will later mark the end of an interior branch.)
It turns out to be convenient for the analysis of branch lengths to consider times relative to the time of observation $T$. We denote the length of the interval between the intermediate time $t$ and the observation time $T$ by $\tau_{e}=T-t$. We further define $\tau=T-\tau_{e}$ as the length of the interval between $\tau_{e}$ and the time of the clade’s origin. Specifically, we denote as $p_{1}^{(\tau_{e})}(\tau)=\tilde{p}^{(T)}_{1}(t=T-\tau_{e})$ the probability that a clade started at time $\tau_{e}+\tau$ in the past, and contains at time $\tau_{e}$ in the past only a single individual that could have extant descendants at the time of observation. Again, setting $\tau_{e}=0$ results in $p_{1}^{(\tau_{e}=0)}(\tau)\equiv p_{1}(\tau)$. For the case of $\tau_{e}>0$ the integrating factor becomes $-\int_{0}^{T-\tau_{e}}\operatorname{\mathrm{d}\!}t^{\prime}\left[\mu+\lambda\left(1-2p_{0}(T-t^{\prime})\right)\right]=-\int_{0}^{\tau}\operatorname{\mathrm{d}\!}\tau^{\prime}\left[\mu+\lambda\left(1-2p_{0}(\tau_{e}+\tau^{\prime})\right)\right]=\log\left(\frac{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}})^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{e}+\tau)})^{2}}\right)$ and we obtain
$p_{1}^{(\tau_{e})}(\tau)=\frac{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}})^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{e}+\tau)})^{2}}\ .$ (7)
This probability differs from $p_{1}(\tau)\equiv p^{(\tau_{e}=0)}_{1}(\tau)$, as for $\tau_{e}>0$ additional lineages have more time to die out than in pendant branches (specifically $\tau_{e}$ more). This will lead to a difference between the branch length statistics of interior and pendant branches. In the following sections, we will measure times $\tau$ relative to the time of observation, with $\tau$ increasing as one looks further into the evolutionary past, see Figure 1.
Figure 1: Times of events on a birth-death tree. Sketch of the situation characterized by $p_{1}^{(\tau_{e})}(\tau)=\tilde{p}_{1}^{(T)}(t)$. From the perspective of a clock starting at the origin of a clade, $p_{1}^{(\tau_{e})}(\tau)=\tilde{p}_{1}^{(T)}(t)$ describes the probability that out of all individuals at time $t$, only one will have extant descendants (depicted in blue) at time $T>t$ when the population is observed. The red lines show lineages which die out before $T$. For convenience, the time of observation is used as the origin of time $\tau$, with time increasing as we look into the evolutionary past. $\tau_{e}$ denotes time point beyond only a single individual has a clade extant at the time of observation, and $\tau$ denotes the time the population has had to grow before $\tau_{e}$. In the corresponding reconstructed phylogenetic tree, there are only the black and blue lineages.
III Branch length distribution: pendant branches
The results from section II can be used directly to determine the statistics of branch lengths of a phylogenetic tree. In this section, we apply them to the so-called pendant branches, which end in a leaf of the phylogenetic tree. We use the approach of Paradis Paradis (2016) to determine branch length distributions and write down concrete expressions for the probability densities. This section and the following are concerned with complete sampling (where the phylogenetic tree is constructed from the entire population alive at a particular time since the start of the birth-death process), in Section VII we will turn to the case where only a finite fraction of the population is available to reconstruct a tree from.
A pendant branch (a branch terminating at the time of observation) of length $\tau$ was created in a birth event at time $\tau$ , where (i) one of the two offspring produces a clade consisting of a single extant member at the time of observation, and (ii) the other offspring produces a clade that did not die out by the time of observation. Figure 2A shows a simple example. Conditions (i) and (ii) are statistically independent events, whose probabilities can be expressed in terms of the quantities computed above. This leaves the rate at which birth events occur per small time interval, which is given by the birth rate $\lambda$ times the population size at time $\tau$. Since the population size is unknown, we sum over all possible population sizes $n$ and their probabilities, yielding that the probability of a birth event at time $\tau$ as the expectation value of the population size multiplied with the birth rate $\lambda$. Plugging $t=T-\tau$ in equation (5) gives the expected population size at time $\tau$ before sampling is proportional to $\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}$.
The relative number of pendant branches of length $\tau$ per small time interval is thus given by
$\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}\left[p_{1}(\tau)p_{>1}(\tau)+p_{>1}(\tau)p_{1}(\tau)+2p_{1}(\tau)^{2}\right]$ (8)
$=2\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}p_{1}(\tau)\left(1-p_{0}(\tau)\right).$ (9)
The factor of two at the end of the first term accounts for the emergence of two branches of length $\tau$ when both offspring only have a single surviving descendant. $p_{>1}(\tau)=\sum_{n=2}^{\infty}p_{n}(\tau)$ denotes the probability that a clade founded at time $\tau$ has more than one extant member at the time of observation. Using equations (2) and (3), one obtains the probability density function of the lengths $\tau$ of pendant branches as
$P_{\text{pend}}(\tau)$ $=2\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}p_{1}(\tau)\left(1-p_{0}(\tau)\right)$ (10)
$=2\lambda\frac{(\lambda-\mu)^{3}\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}}{\left(\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}-\mu\right)^{3}}\ .$
For a finite time since the start of the birth-death process, $\tau$ has a finite support and hence $P_{\text{pend}}(\tau)$ cuts off to zero when $\tau$ is larger than the time $T$ since the start of the birth-death process. The normalizing factor of (10) is then found to be $(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T})^{2}/(\lambda(\lambda-(\lambda-2\mu)\operatorname{\mathrm{e}}^{-2(\lambda-\mu)T}-2\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T}))$. We focus on the particularly simple case of an asymptotically long running time $T\to\infty$. In this case, the cutoff moves to infinity and we find the normalizing factor in (10) to be one, making (10) a proper probability density function. We will focus on the asymptotic case $T\to\infty$ also when treating branches in the tree’s interior.
Curiously, distribution (10) coincides with the distribution of pendant branches conditioned on the population size $N$ (statistical scenario ii)), derived by Stadler and Steel Stadler and Steel (2012) as well as Mooers et al. Mooers et al. (2012). For details, see Appendix .1.
IV Branch length distribution: interior branches above pendant branches
This approach to pendant branches can be extended to the interior branches of a phylogenetic tree. We proceed to go up the phylogenetic tree and first consider branches one level up from pendant branches. We call branches that end in an internal node that is the beginning of a pendant branch level-two branches. Correspondingly, pendant branches are level-one branches. Figure 2B shows the situation: a pair of successive level-two and level-one branches of lengths $\tau_{2}$ and $\tau_{1}$, respectively, arose in a birth event at time $\tau_{2}+\tau_{1}$ in the past, where one of the two offspring has extant descendants (d-f), and the other has no extant descendants apart from the offspring of individuals arising in another birth event at time $\tau_{1}$, which gives rise to two offspring. Again, one of these has at least one extant descendant (b-c), and the other has exactly one (a), establishing the pendant branch of length $\tau_{1}$, see Fig. 2B.
ARefer to caption                    BRefer to caption
CRefer to caption
DRefer to caption      ERefer to caption
Figure 2: Statistics of pendant and interior branches. (A) This schematic phylogenetic tree shows birth events (circles) and death events ($\times$). The pendant branch of length $\tau$ on the left arose from a birth event in the past (solid circle), where one offspring (shown left) has exactly one extant descendant, and the other offspring (shown right) has at least one extant descendant. This particular pendant branch has one generation between its origin at time $\tau$ measured from the time of observation in the past and the present (not counting the birth event at $\tau$). (B) A birth event (top solid node) leading to a level-two branch of length $\tau_{2}$ and a level-one (pendant) branch of length $\tau_{1}$. (C) A schematic phylogenetic tree with branches of different levels, see text. (D) The distribution of branch lengths $\tau_{2}$ of second-level branches in the tree of a simulated population grown from a single individual over time $T=25$ as described in the text (grey bars), compared to the theoretical probability density function (11) in the asymptotic case of large $T$ marginalized over the pendant branch lengths $\tau_{1}$ (blue line). (E) The distribution of interior branch lengths in the same tree as before (grey bars), compared to the theoretical density function (14) (blue line).
Since the probability of a birth event at time $\tau_{1}+\tau_{2}$ was $\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{1}+\tau_{2})}$ (up to a normalizing prefactor, see below), the relative number of level-two and level-one branches of lengths $\tau_{2}$ and $\tau_{1}$ per small time intervals is thus
$P_{\text{level}}(\tau_{1},\tau_{2})=2\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{1}+\tau_{2})}p_{1}(\tau_{1})\left(1-p_{0}(\tau_{1})\right)2\lambda p^{(\tau_{1})}_{1}(\tau_{2})\left(1-p_{0}(\tau_{1}+\tau_{2})\right)\ .$ (11)
Here $p^{(\tau_{1})}_{1}(\tau_{2})$ denotes the probability that all birth events taking place along the level-two branch of length $\tau_{2}$ lead to additional lineages that die out by the time the population is observed. More precisely, it is the probability that a given individual alive at time $\tau_{1}+\tau_{2}$ has exactly one descendant alive at time $\tau_{1}$ that will have extant descendants. It is given by equation (7).
Equation (11) gives the joint probability density function of branch lengths of pendant and level-two branches up to a normalization factor, which again turns out to be one. Figure 2D compares this result to numerical simulations. A population with birth rate $\lambda=1$ and death rate $\mu=0.75$ was grown from a single individual until time $T=25$ (where the population contained 3567 individuals), and the phylogenetic tree was reconstructed. The lengths of level-two branches were collected in a histogram and compared to the joint probability $P_{\text{level}}(\tau_{1},\tau_{2})$ given by (11) marginalized with respect to $\tau_{1}$.
This approach can be applied to higher levels in the phylogenetic tree. The joint probability density function of branch lengths $\tau_{1},\tau_{2}$ and $\tau_{3}$ in a lineage with level-one, -two, and -three branch is
$P_{\text{level}}(\tau_{1},\tau_{2},\tau_{3})=2\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{1}+\tau_{2}+\tau_{3})}p_{1}(\tau_{1})\left(1-p_{0}(\tau_{1})\right)2\lambda p^{(\tau_{1})}_{1}(\tau_{2})\left(1-p_{0}(\tau_{1}+\tau_{2})\right)2\lambda p^{(\tau_{1}+\tau_{2})}_{1}(\tau_{3})\left(1-p_{0}(\tau_{1}+\tau_{2}+\tau_{3})\right)\ .$ (12)
These densities describe the statistics of branches as one goes up the tree from the different leaves; starting from one particular leaf gives one set of branch lengths $\tau_{1},\tau_{2},\ldots$, and starting from another leaf gives another set of branch lengths. However, at higher levels than the first, not all branches can be uniquely assigned to a specific level. For instance, the branch marked $\ast$ in Figure 2C contributes both to the distribution of $\tau_{2}$ (going up the tree from the leaf $a$) and to the distribution of $\tau_{3}$ (going up from the leaves marked $b$ or $c$).
V Branch length distribution: interior branches
As a next step, we derive the distribution of internal branch lengths, i.e., all branches that are not pendant. (We exclude the branch between the start of the birth-death process to the first bifurcation.) To this end, we calculate the joint probability of an interior branch starting at time $\tau_{s}$ and ending at $\tau_{e}<\tau_{s}$. As before, the probability per unit time of a birth event occurring at time $\tau_{s}$ is proportional to $\lambda\operatorname{\mathrm{e}}^{-(1-q)\tau_{s}}$. Focusing for now on only one of the two clades that started at $\tau_{s}$ (called $A$), the probability of it running until $\tau_{e}$ with only one extant descendant at the time of observation is $p_{1}^{(\tau_{e})}(\tau_{s}-\tau_{e})$. The probability per small time interval of the branch ending at $\tau_{e}$ is equal to the probability per small time interval of a birth event at that time where none of the two emerging clades dies out. This is given by $\lambda(1-p_{0}(\tau_{e}))^{2}$.
Two cases must be considered for the other clade (called $B$) emerging at $\tau_{s}$. With probability $p_{1}(\tau_{s})$ this clade becomes a pendant branch. In this case, a factor of two arises due to symmetry (either clade can be called $A$). An alternative scenario is the clade having more than one extant descendant. This happens with probability $p_{>1}(\tau_{s})=1-p_{0}(\tau_{s})-p_{1}(\tau_{s})$. In this case, the clade also has an interior branch starting at $\tau_{s}$ contributing to the branch length distribution. A factor of two accounts for this.
Putting these steps together, the relative probability of an interior branch running from $\tau_{s}$ to $\tau_{e}$ (up to a normalization factor found to be one) is given by
$P_{\text{int}}(\tau_{s},\tau_{e})$ $=\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{s}}p_{1}^{(\tau_{e})}(\tau_{s}-\tau_{e})\lambda\left(1-p_{0}(\tau_{e})\right)^{2}\left[2p_{1}(\tau_{s})+2p_{>1}(\tau_{s})\right]$
$=2\lambda^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{s}}p_{1}^{(\tau_{e})}(\tau_{s}-\tau_{e})\left(1-p_{0}(\tau_{e})\right)^{2}\left(1-p_{0}(\tau_{s})\right)$ (13)
$=2\lambda^{2}(\lambda-\mu)^{3}\frac{\operatorname{\mathrm{e}}^{(\lambda-\mu)(\tau_{s}+\tau_{e})}}{(\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau_{s}}-\mu)^{3}}\ .$
Marginalizing over $\tau_{s}$ for $\tau=\tau_{s}-\tau_{e}$ results in the density of interior branch lengths in the entire tree
$P_{\text{int}}(\tau)$ $=(\lambda-\mu)^{2}\frac{2\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{\left(\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}-\mu\right)^{2}}.$ (14)
Figure 2E compares this result to the empirical distribution of a tree reconstructed from a population of age $T=25$ grown from a single individual with birth rate $\lambda=1$ and death rate $\mu=0.75$ as described above.
So far, we have again only considered the asymptotic case of $T\rightarrow\infty$. For finite $T$, the probability of $\tau_{s}>T$ is 00, resulting in a normalization factor of $(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T})^{2}/(a(1-\operatorname{\mathrm{e}}^{-(\lambda-\mu)T}))^{2}$ for (13).
VI Number of generations along branches
Birth events where one of the two newly founded clades dies out do not lead to bifurcations in the reconstructed tree. Such birth events are not visible from the reconstructed tree alone, yet they contribute to the number of generations along a lineage. In this section, we discuss the statistics of the number of generations along a particular branch of a phylogenetic tree.
The mean number $i$ of generations along a particular branch starting at time $\tau_{s}$ and ending at time $\tau_{e}$ (not counting the founding event of the branch) has been calculated by Bokma, van den Brink, and Stadler Bokma et al. (2012) as
$\langle i\rangle(\tau_{s},\tau_{e})=2\lambda\int_{\tau_{e}}^{\tau_{s}}\operatorname{\mathrm{d}\!}\tau\,p_{0}(\tau)=2\left[\mu\tau-\log\left(\lambda-\mu e^{-(\lambda-\mu)\tau}\right)\right]_{\tau_{e}}^{\tau_{s}}$ (15)
based on birth events occurring along a branch at rate $\lambda$ and the probability $p_{0}(\tau)$ that one given additional lineage has died out by time $\tau$ given by (2). The factor $2$ arises because either of the two offspring can be the one to die out. The number of generations is Poisson distributed (and hence fully characterized by the mean (15)), because successive birth events are statistically independent Bokma et al. (2012).
Together with the densities of branch lengths from the previous sections, we can now compute the probability distribution for the number of generations along a randomly chosen branch. In the case of pendant branches and asymptotically old trees, it is given by
$P_{\text{pend}}(i)=\int_{0}^{\infty}\operatorname{\mathrm{d}\!}\tau\,P_{\text{pend}}(\tau)\operatorname{\mathrm{e}}^{-\langle i\rangle(\tau,0)}\frac{\left[\langle i\rangle(\tau,0)\right]^{i}}{i!}\ .$ (16)
Similarly, the probability of finding $i$ generations along a randomly chosen interior branch is
$P_{\text{int}}(i)=\int_{0}^{\infty}\operatorname{\mathrm{d}\!}\tau_{s}\int_{0}^{\tau_{s}}\operatorname{\mathrm{d}\!}\tau_{e}\,P_{\text{int}}(\tau_{s},\tau_{e})\operatorname{\mathrm{e}}^{-\langle i\rangle(\tau_{s},\tau_{e})}\frac{\left[\langle i\rangle(\tau_{s},\tau_{e})\right]^{i}}{i!}\ .$ (17)
VII Finite sampling
Usually only a subset of individuals sampled from the population is available to construct the phylogenetic tree from and this sampling affects the statistics of branch lengths. To model the process of sampling from the extant population we follow the standard approach and choose extant individuals at the time of observation independently and with equal probability $\rho$ Yang and Rannala (1997); Stadler (2010); Stadler and Steel (2012). The preceding results on branch length distributions can be easily generalized to a sampling probability $\rho<1$ using the probabilities that a clade of a particular age has no (or one) member which is sampled at a particular time.
Yang and Rannala Yang and Rannala (1997) have calculated the probabilities that a birth-death process starting with a single individual at time zero has no sampled offspring or a single sampled offspring, respectively, as
$p_{0}^{(\rho)}(\tau)$ $=1-\frac{\rho(\lambda-\mu)}{\rho\lambda+\left((1-\rho)\lambda-\mu\right)\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}$ (18)
$p_{1}^{(\rho)}(\tau)$ $=\frac{\rho(\lambda-\mu)^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{(\rho\lambda+\left((1-\rho)\lambda-\mu\right)\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau})^{2}}\ .$ (19)
[…]
$p_{1}^{(\rho)(\tau_{e})}(\tau)$ $=\frac{\left(\rho\lambda-(\mu-(1-\rho)\lambda)\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}}\right)^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{\left(\rho\lambda-(\mu-(1-\rho)\lambda)\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{e}+\tau)}\right)^{2}}\ .$ (20)
Note that in this case, $p_{1}^{(\rho)}(\tau)=\rho p_{1}^{(\rho)(\tau_{e}=0)}(\tau)$, because we need to include the probability that the lineage running until $\tau_{e}=0$ is actually sampled.
To derive the probability density function of branch lengths under sampling, these probabilities can be used directly in the densities of pendant and interior branch lengths at full sampling ($\rho=1$), (10)-(13), which depend only on $p_{0}$, $p_{1}$, and the relative population sizes at different times. In particular, we obtain
$P^{(\rho)}_{\text{pend}}(\tau)$ $=2\lambda\rho(\lambda-\mu)^{3}\frac{\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}}{\left(\lambda\rho\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}+\lambda(1-\rho)-\mu\right)^{3}}$ (21)
$P_{\text{int}}^{(\rho)}(\tau_{s},\tau_{e})$ $=2(\lambda\rho)^{2}(\lambda-\mu)^{3}\frac{\operatorname{\mathrm{e}}^{(\lambda-\mu)(\tau_{s}+\tau_{e})}}{\left(\lambda\rho\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau_{s}}+\lambda(1-\rho)-\mu\right)^{3}}$ (22)
$P_{\text{int}}^{(\rho)}(\tau)$ $=(\lambda-\mu)^{2}\frac{2\lambda\rho-\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}\left(\mu-\lambda(1-\rho)\right)}{\left(\lambda(1-\rho)-\mu+\lambda\rho\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}\right)^{2}}$ (23)
A normalizing factor of $\rho^{-1}$ applies to all three densities and is included in the expressions above. The result for pendant branches coincides with the result of Stadler and Steel Stadler and Steel (2012) at fixed number of samples. However, their result was restricted to a regime of high death rate or high sampling probability $1\geq\mu/\lambda\geq 1-\rho$. The density function (21) also holds outside these limits.
Stadler Stadler (2009) found that the probability distribution of bifurcation times is not unique to a particular set of parameters, but that a birth-death process with birth rate $\lambda$, death rate $\mu$ and sampling probability $\rho$ yields the same distribution as a process with parameters $\rho^{\prime}$, $\lambda^{\prime}=\rho\lambda/\rho^{\prime}$, $\mu^{\prime}=\mu-\lambda(1-\rho/\rho^{\prime})$. We find the same for the distributions of branch lengths (21)-(23), as $\lambda^{\prime}-\mu^{\prime}=\lambda-\mu$, $\rho^{\prime}\lambda^{\prime}=\rho\lambda$ and $\lambda^{\prime}(1-\rho^{\prime})-\mu^{\prime}=\lambda(1-\rho)-\mu$ (which is the same argument used in Stadler (2009, 2013a)). Therefore, inference of both growth parameters and the sampling probability from the distribution of branch lengths is impossible. However, an upper bound on the sampling probability can be determined: Given a branch distribution generated with (unknown) parameters $(\lambda,\mu,\rho)$, the parameter space of all sets of $(\lambda^{\prime},\mu^{\prime},\rho^{\prime})$ which lead to the same distribution is restricted to $\mu^{\prime}\geq 0$ and $0<\rho\leq 1$. With the transformation rules above, it is easy to see that these inequalities are only fulfilled if $0<\rho^{\prime}\leq\min(\rho\lambda/(\lambda-\mu),1)$. This means that for branch distributions generated with birth rate $\lambda$, death rate $\mu$ and sampling probability $\rho<1-\mu/\lambda$, there exists a maximal sampling probability $\rho^{\prime}<1$ for the family of possible parameter sets.
The lengths of pendant and interior branches behave very differently under sampling. Pendant and interior branch lengths are compared in Figure 3A. The rows correspond to different sampling probabilities $\rho$, the first column shows pendant, the second column interior branches. Histograms show the branch length distributions of $100$ simulations of trees, while the blue lines depict the analytic results. Pendant branches in the first column get longer with decreasing sampling probability $\rho$. On the other hand, interior branches quickly reach an asymptotic distribution as $\rho$ is decreased and then do not get any longer as $\rho$ is decreased further, see the second column of Figure 3A. The red dashed line depicts the asymptotic state, which can be derived as
$P^{(\rho\rightarrow 0)}_{\text{int}}(\tau)=\lim_{\rho\rightarrow 0}P^{(\rho)}_{\text{int}}(\tau)=(\lambda-\mu)\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}$ (24)
and is thus an exponential distribution with parameter $\lambda-\mu$.
ARefer to caption BRefer to caption                   C Refer to caption
Figure 3: Pendant and interior branches at different sampling probabilities $\rho$. (A) We compare the analytical results for the probability density of the length of pendant and interior branches under sampling ((21) and (23) respectively, solid blue lines) to numerical simulations. For the latter, we show the distribution of branches from 100 trees. For every tree, we simulated the growth of a population with $\lambda=1$ and $\mu=0.75$ over time $T=40$, restarting when the population died out or the final population size was not within $2\times 10^{4}$ and $3\times 10^{5}$ (grey bars). The first row of plots shows the branch length distribution of the full trees. For the next row of plots, individuals were sampled with probability $\rho=0.1$ and the corresponding trees were reconstructed. For the third row, each of the previous samples was sampled again with probability $0.1$, corresponding to an overall sampling probability of $\rho=0.01$. The same was repeated for the forth row of plots, yielding $\rho=0.001$. Every tree contributes equally to the empirical distributions depicted. The first column shows the distributions of pendant branch lengths, which shift to longer branches as $\rho$ decreases. The second column of plots shows the distribution of interior branches, quickly reaching the asymptotic distribution (24) (red dashed line) as $\rho$ becomes small. (B) The distribution of the number of generations along interior branches in the asymptotic case of small sampling probabilities (25). Different colours represent different relative death rates $\mu/\lambda$. (C) A toy example of how sampling affects the length of pendant branches, see text.
The difference between pendant and interior branches under sampling has a simple qualitative explanation, which we explore within a toy problem. Consider a reconstructed tree which is perfectly balanced; both clades defined by each bifurcation have the same size. An example is shown in Fig. 3C. Then a fraction $\rho$ of the extant nodes is sampled. Under sampling, a pair of pendant branches linked to a shared ancestor is defined by the birth of a clade (shown in blue in Fig. 3C) whose size $n$ at the time of observation is defined by $n\rho\approx 2$ (so two of its members are expected to be sampled). The number of branch levels between this founding birth event and the time of observation is $\log_{2}(n)$, so the pendant branch lengths are increased under sampling by a factor of $1-\log_{2}\rho\approx-\log_{2}\rho$ relative to the fully sampled case. (This argument assumes that on the fully sampled reconstructed tree, pendant and interior branches are approximately of the same length.) On the other hand, branches that are interior in the reconstructed tree (i.e. lie above birth event shown in blue in Fig. 3C) have their lengths unchanged by sampling. The distribution of interior branches only changes because under sampling, fewer and older branches are counted as interior. Older branches tend to be a little longer (see equation (13)), because any additional lineages that arose along them had more time to die out until the time of observation.
Following Section VI and equations (21)-(23), we also compute the distribution of generations along branches of a sampled tree. Crucially, since the mean number of generations (equation (15)) along a branch of length $\tau$ converges to $2\mu\tau$ as the branch moves further away from the time of observation (as interior branches do when taking the limit $\rho\rightarrow 0$), we obtain the probability of finding $i$ generations along an interior branch in this limit
$P^{(\rho\rightarrow 0)}_{\text{int}}(i)=\int_{0}^{\infty}\operatorname{\mathrm{d}\!}\tau\,P^{(\rho\rightarrow 0)}_{\text{int}}(\tau)\operatorname{\mathrm{e}}^{-2\mu\tau}\frac{(2\mu\tau)^{i}}{i!}=(2\mu)^{i}(\lambda-\mu)(\lambda+\mu)^{-i-1}\ .$ (25)
This asymptotic result is depicted for different relative death rates $\mu/\lambda$ in Figure 3B.
VIII Branch length distributions conditioned on the population size at the time of observation
We now turn to the second statistical scenario, where the birth-death process is conditioned on the number $N$ of extant individuals rather than a particular time of observation (scenario ii) above). This case is harder to treat than scenario i), because with a constraint on the final number of individuals, the dynamics of different clades are in general correlated. We set up a framework to calculate branch length distributions based on tracking the dynamics of the birth-death process first backwards in time from the time of observation to the emergence of a particular branch and then again forward in time to the time of observation. For pendant branches, this approach recovers the results of Stadler, Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012), but it also allows treating interior branches.
We first illustrate the method with a toy problem: We calculate the distribution of the length of pendant branches when the number of individuals at the time of observation is $N=2$. A key tool is the probability that a birth-death process with birth rate $\lambda$ and death rate $\mu$, starting with $n$ individuals at a given time has $N$ individuals a time interval $t$ later, denoted $p_{N,n}^{\lambda,\mu}(t)$.
Figure 4: Pendant branch formation for populations of size $N=2$: The red branch marks the pendant branch arising from a birth event at time $\tau$ in the past. All other $n-1$ individuals that were alive at time $\tau$ need to die out before the population is observed, while the second descendant from the birth event needs to survive with clade size one. The arrows on the right illustrate the conditional probabilities of having $N$ extant descendants given that $n$ individuals at time $\tau$ before observation (forwards process) and the probability that given $N$ observed individuals there were $n$ individuals at time $\tau$ before observation (backwards process).
First we go backwards in time from the time of observation to the time $\tau$. Conditioned on the number of individuals at the time of observation being $N$, the statistics of the number of individuals before the time of observation is described by a birth-death process with time running backwards and with the rates of birth and death exchanged, so the rate of birth is $\mu$ and the rate of death is $\lambda$ Ignatieva et al. (2020). The probability of having $n$ individuals at time $\tau$ before the time of observation is thus $p_{n,N}^{\mu,\lambda}(\tau)$. The number of individuals $n$ will eventually be summed over.
With $n$ individuals at time $\tau$, the expected number of birth events at time $\tau$ (per small time interval) leading to pendant branches, conditioned on $N=2$ individuals at the time of observation is
$\lambda np_{1}^{2}(\tau)p_{0}^{n-1}(\tau)/p_{2,n}^{\lambda,\mu}(\tau)\ .$ (26)
$\lambda n$ gives the probability per small time interval that one of the $n$ individuals divides during this small interval, $p_{1}^{2}(\tau)$ is the probability that the two resulting offspring each have one extant offspring at the time of observation, leading to pendant branches of length $\tau$. $p_{0}^{n-1}(\tau)$ is the probability that all other $n-1$ individuals die out by the time of observation, leading to a population of size $N=2$. The denominator arises from the conditioning on $N=2$ individuals. What we have done is first asking about the statistics of the number of individuals at a certain time $\tau$ in the past, given $N$ individuals at some time of observation, and then asking about the probability that a tree with a particular feature (here a pendant branch that emerged at time $\tau$) arises from those forward-time trajectories ending in $N$ individuals at the time of observation. For an illustration, see Figure 4.
Up to a normalizing constant, the probability density function of the pendant branch lengths is thus
$\sum_{n=1}^{\infty}\lambda np_{1}^{2}(\tau)p_{0}^{n-1}(\tau)\frac{p_{n,2}^{\mu,\lambda}(\tau)}{p_{2,n}^{\lambda,\mu}(\tau)}\ .$ (27)
In Appendix .1 we show how this expression evaluates to $2\lambda\frac{p_{1}^{2}}{1-p_{0}}=2\lambda\frac{(\lambda-\mu)^{3}\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}}{\left(\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}-\mu\right)^{3}}$ and a normalisation constant of one. We find the same result for a final population size $N\geq 2$. This reproduces the result by Stadler and collaborators on the distribution of pendant branches conditioned on the size $N$ of the population at the time of observation. Curiously, this distribution also coincides with the distribution of pendant branch lengths under scenario i) in the limit of large $\tau$ (equation (10)).
The same approach can also be used to calculate the distribution of interior branch lengths conditioned on the population size. Again we use the birth-death process with time running backwards to compute the probability that at a time $\tau_{s}$ in the past there were $n$ individuals, one of which divided establishing a new branch. Running time forwards again and conditioning on the final population size $N$, we calculate the probability that the new branch ends at $\tau_{e}$ in another birth event, with both descendants having extant offspring. The conditions that lead to a branch of a particular length are the same manner as in scenario i), the difference is the conditioning on the final population size in two directions of time. The calculation is given in Appendix .2.
The resulting distribution of interior branch lengths under scenario ii) depends on $N$. However, the dependency quickly vanishes with $\left(\frac{\lambda}{\mu}p_{0}(\tau_{s})\right)^{N-1}$ when increasing the population size. The asymptotic distribution for large $N$ coincides with the interior branch length distribution under scenario i) in the limit of large times, see Appendix .2.
IX Discussion
We have calculated the statistics of branch lengths (in calendar time) of phylogenetic trees generated by a constant-rate birth-death process and reconstructed from the population at a single moment in time, with a particular focus on how branch lengths change when the population is sampled at different sampling probabilities. We look at two distinct statistical scenarios: sampling at a particular time since the birth-death process started and sampling when the population has a given size $N$.
For the first scenario, we generalize the results of Paradis Paradis (2016) to a finite sampling probability. We find that pendant branches and interior branches behave very differently under sampling: only pendant branch lengths increase indefinitely as the sampling probability is decreased to zero, whereas the distribution of interior branches quickly reaches an asymptotic. This behaviour allows placing bounds on the model parameters (birth and death rates, sampling probability) on the basis of empirical trees. A second consequence of this behaviour under sampling concerns the number of generations along interior branches: even at low sampling probabilities, the number of generations along interior branches can be small — depending on the ratio of death and birth rates, but asymptotically independently of the sampling. Under cellular reproduction, when mutations occur at birth events, this means that the number of mutations along a branch can be shaped by the (few) birth events along that branch rather than accumulating continuously over time. The consequences and applications of this result will be explored in a future publication.
For the second scenario, a fixed given final population size $N$, we develop a statistical framework based on first going backwards in time from a fixed population size, and then forwards again while conditioning on that population size. The latter step entails calculating the probability of a population with a given final size and a particular feature of its phylogenetic tree (like a branch of a given length), relative to the probability of a population with a certain size. The combination of going backwards and forwards in time is similar in spirit to the celebrated Keldysh––Schwinger formalism in non-equilibrium quantum physics Kamenev (2011). Our approach simplifies the derivation of the distribution of pendant branch lengths by Stadler and Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012) and allows also to treat the length distribution of interior branches.
The interior branch distributions of the asymptotic cases of both scenarios (large times and large population sizes) turn out to coincide. In case of pendant branches, the asymptotic distribution in the limit of large times since the start of the birth-death process coincides with the distribution in trees conditioned on population size N, even when N is not large, as the latter distribution is independent of $N$ Stadler and Steel (2012).
Acknowledgements.
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) grant SFB1310/2 - 325931972. Many thanks to Tibor Antal, Samuel Johnston, Joachim Krug for discussions, and to Tanja Stadler for comments throughout and especially for helpful advice on section VII. We thank the computing center of the University of Cologne RRZK for computing time on the DFG-funded CHEOPS cluster (grant number INST 216/512/1FUGG).